OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n. The integer n^3 + 2n is divisible by 3 for every positive integer n.

OpenStudy (anonymous):

just for clarification...thats \(n^3+2n\) :)

OpenStudy (anonymous):

@mukushla yeah sorry!

OpenStudy (anonymous):

Well, consider \(n^3+2n=n(n^2+2)\). If \(3\,|\,n\) it's obvious. If \(3\not|\,n\) we have \(n\equiv1\text{ or }n\equiv 2\pmod3\) hence \(n^2\equiv1\pmod3\) and thus \(n^2+2\equiv0\mod3\)... so we're done :-) sorry this isn't an inductive proof but I wanted to play with modular arithmetic. Anyways, where were we? Let's assume our statement is true for some positive integer \(k\), i.e. \(k^3+2k\) is divisible by \(3\). Anyways, consider then \(k+1\) and we have:$$(k+1)^3+2(k+1)=k^3+3k^2+3k+1+2k+2=(k^3+2k)+3(k^2+k+1)$$; clearly since \(3\) divides \(k^3+2k,3(k^2+k+1)\) it follows that \(3\) divides \((k+1)^3+2(k+1)\). Now, we merely show it starts for the first positive integer \(n=1\) and we've proved it all for all integers:$$1^3+2(1)=1+2=3$$and \(3\) divides \(3\). It follows that \(n^3+2n\) is divisible by \(3\) for all positive integers \(n\)

OpenStudy (anonymous):

Thanks! I wasn't sure where to go after k^3+3k^2+5k+3 c:

OpenStudy (anonymous):

@Adri1580 generally since you're assuming \(k^3+2k\) is divisible by \(3\) you want to those terms 'out' and consider the remaining ones to show it holds for \((k+1)^3+2(k+1)\) as well :-)

OpenStudy (anonymous):

Thanks again for everything :)

OpenStudy (anonymous):

move* those terms 'out' and to the side so you can focus on the new remaining ones**